The comparative balance sheets of Sloan Company reveal that accounts receivable (before deducting allowances) increased by $15,000 in 2016. During the same time period, the allowance for uncollectible accounts increased by $2,100. If sales revenue was $120,000 in 2016 and bad debts expense was 1% of sales, how much cash was collected from customers during the year?  
Answer #1

Bad debt expense = 120000*1% = 1200

Cash Collection = Sales revenue-Account receivable increase

= 120000-15000

Cash collection = 105000